OpenStudy (anonymous):

cos(5°) sin(-4°) - cos(4°) sin(5°) = sin( )

OpenStudy (solomonzelman):

sin(-4)=-sin(4) cos(5°) sin(-4°) - cos(4°) sin(5°) = -cos(5°) sin(4°) - cos(4°) sin(5°) = -(cos(5°) sin(4°) + cos(4°) sin(5°) )= you got it from here, use sin(a+b)=sin(a)cos(b)+sin(b)cos(a)

OpenStudy (anonymous):

9?

OpenStudy (solomonzelman):

well, cos(5°) sin(4°) + cos(4°) sin(5°) = sin(9) So, .... -(cos(5°) sin(4°) + cos(4°) sin(5°) )= -sin(9)

OpenStudy (anonymous):

or will it be -5+4 so -1

OpenStudy (anonymous):

so -9

OpenStudy (solomonzelman):

sin of -9, not just -9.